# FedLoDrop: Federated LoRA with Dropout for Generalized LLM Fine-tuning

### v1

@@ II-B LoRA Dropout
The original dropout technique was proposed in [32] to avoid overfitting during training. In standard dropout, each neuron in the network is dropped from the network with a certain probability. Given that dropout techniques have proven effective in controlling overfitting, we introduce a LoRA dropout framework to enhance generalization when adapting to downstream tasks. The loss function can be re-written as $\mathcal{L}=\frac{1}{|\mathcal{D}|}\sum_{i\in\mathcal{D}}\ell\left(\Delta\bm{\theta}(\bm{m}_{t});\bm{\theta}_{0},{x}_{i}\right),$ where $\Delta{\bm{\theta}}(\bm{m}_{t})$ is the LoRA parameters after the dropout, $\bm{m}_{t}$ is the concatenation of masks of all LoRA modules, which varies in different training rounds, and $\bm{\theta}_{0}$ is the original parameters of the pre-trained model.
@@ II-B1 Forward Pass
After dropout, the output of the $u$-th layer is given by
$\hat{\bm{h}}_{u,i,t}$ $=(\bm{W}_{u,0}+\hat{\bm{B}}_{u,t-1}\hat{\bm{A}}_{u,t-1})\bm{f}_{u-1,i,t}.$ (12)
@@ II-C1 Procedure and Algorithm
In this case, each device $k$ holds a local dataset, denoted by $\mathcal{D}_{k}=\{\bm{x}_{i}|i=1,2,\ldots,|\mathcal{D}_{k}|\}$, where $\bm{x}_{i}$ is the $i$-th data sample and the size of $\mathcal{D}_{k}$ is $|\mathcal{D}_{k}|$. Denote $\Delta{\bm{\theta}}(\bm{m}_{k,t})$ as the LoRA parameters after the dropout for the $k$-th device, $\bm{m}_{k,t}$ as the concatenation of all dropout masks of LoRA modules, which varies across different training rounds and different devices, and $\bm{\theta}_{0}$ as the original parameters of the pre-trained model, which has been stored in each device in advance. The objective is to minimize the global loss function, given by $\mathcal{L}=\sum_{k=1}^{K}\frac{\left|\mathcal{D}_{k}\right|}{|\mathcal{D}|}\ell_{k}\left(\Delta{\bm{\theta}}(\bm{m}_{k,t});\bm{\theta}_{0},\mathcal{D}_{k}\right),$ where $\mathcal{D}=\{\mathcal{D}_{k}\}$ is the global dataset, the local loss function of the $k$-th device can be written as $\ell_{k}\left(\Delta{\bm{\theta}}(\bm{m}_{k,t});\bm{\theta}_{0},\mathcal{D}_{k}\right)=\frac{1}{|\mathcal{D}_{k}|}\sum_{{x}_{i}\in\mathcal{D}_{k}}\ell_{k,i}\left(\Delta{\bm{\theta}}(\bm{m}_{k,t});\bm{\theta}_{0},\bm{x}_{i}\right).$ Similar to (11), FedLoDrop can be written as
@@
  • LoRA Sub-adapter Generation: The server adopts the LoRA dropout technique introduced in Section II-B to generate a sub-adapter for each device.
@@ III-B Generalization Error Analysis
In (18), $\ell\left(\bm{\theta},\bm{x}_{j}\right)$ represents the loss for sample $\bm{x}_{j}$ given model parameters $\bm{\theta}$. According to [29], if the following requirements are met: $\ell$ is $\eta$-Lipschitz, $\bm{\theta}_{\ell_{k,\lambda}}(\mathcal{D}_{k})$ is close to $\bm{\theta}_{\ell_{k,\lambda_{t}}}(\mathcal{D}_{k}^{j})$, Hessian matrix $\nabla^{2}{\ell_{k,\lambda_{t}}}(\bm{\theta}_{\ell_{k,\lambda_{t}}}(\mathcal{D}_{k}))$ at $\bm{\theta}_{\ell_{k,\lambda_{t}}}(\mathcal{D}_{k})$ is positive-semidefinite with a singular value decomposition $U_{k,t}\text{diag}(\Lambda_{k,t})U_{k,t}^{-1}$, and $\Lambda_{k,t,\text{min}}=\text{min}\{\Lambda_{k,t,1},\cdots,\Lambda_{k,t,m}\}$, then the LoRA dropout algorithm optimizing $\ell_{k,\lambda_{t}}$ on $|\mathcal{D}_{k}|$ has an upper bound of PHS of  [^2] We only require the regularization coefficient to be sufficiently large (i.e., $\lambda_{t}\geq-\frac{1}{2}\Lambda_{k,t,\min}$), avoiding overly strong assumptions on the objective function. Since uniform stability is stricter and argument stability focuses on parameters, we adopt PHS. Despite the non-convexity of LLM training, which may introduce theoretical gaps, the resulting bound provides meaningful qualitative insights into the client numbers and dropout rates. Tighter bounds under weaker assumptions remain an exciting direction for future work.
@@
Theorem 3 elucidates the connection between generalization error and dropout rate. It is shown that increasing the dropout rate decreases the gap between empirical and generalization errors but concurrently increases empirical error. Therefore, an appropriate dropout rate would help balance a trade-off between overfitting and underfitting. A larger value of $\lambda_{t}$ more strongly enforces the sparsity constraint, resulting in a more compact parameter update and a lower theoretical generalization error bound. This, in turn, reduces the risk of overfitting and enhances the model’s expected performance on unseen data. While the theorem does not explicitly characterize the relationship between generalization error and LoRA rank, it suggests that decreasing the rank-enhancing parameter sparsity-typically correlates with reduced generalization error. This implies that a model with a lower effective rank tends to have better generalization performance [35].
@@ III-C Upper Bounding Loss Descent
$\quad\quad-\alpha_{B}\Delta\hat{\bm{B}}_{u,t-1}\bm{A}_{u,t-1}$
@@ VI-A Effects of Dropout Rate
FedIT with Gaussian Dropout
62.08 77.98 90.20 96.22 94.25 91.72 85.41
FedIT with Gaussian Dropout
61.82 77.98 89.22 95.41 93.85 91.47 84.96
@@ VI-D Effects of Parameters and Models
Fig. 6 presents the effects of parameters on the performance. From Fig. 6(a) and Fig. 6(b), one can see that in both models, a larger number of devices generally results in a lower testing accuracy due to the difficulty in adapting to one global model. As shown in 6(c), extremely increasing the rank results in performance deterioration, likely due to overfitting. This highlights that performance does not always improve with more trainable parameters across different tasks.
@@ A-B PHS upper bound of FedLoDrop
Integrating the result above to Lemma 5 and substituting the regularization coefficient with $\lambda(2\gamma_{k,t}-\gamma_{k,t}^{2})$ finalizes the proof.
@@ Appendix D Proof Of Lemma 4
where, $o_{11}=\frac{4\gamma_{1}^{2}}{(a_{1}-\tilde{\gamma_{1}}^{2})^{3}\sqrt{\frac{1}{a_{1}-\tilde{\gamma_{1}}^{2}}+\frac{1}{a_{2}-\tilde{\gamma_{2}}^{2}}}}-\frac{\gamma_{1}^{2}}{(a_{1}-\tilde{\gamma_{1}}^{2})^{4}(\frac{1}{a_{1}-\tilde{\gamma_{1}}^{2}}+\frac{1}{a_{2}-\tilde{\gamma_{2}}^{2}})^{3/2}}+\frac{1}{(a_{1}-\tilde{\gamma_{1}}^{2})^{2}\sqrt{\frac{1}{a_{1}-\tilde{\gamma_{1}}^{2}}+\frac{1}{a_{2}-\tilde{\gamma_{2}}^{2}}}}$, $o_{12}=o_{21}=-\frac{\gamma_{1}\gamma_{2}}{(a_{1}-\tilde{\gamma_{1}}^{2})^{2}(a_{2}-\tilde{\gamma_{2}}^{2})^{2}(\frac{1}{a_{1}-\tilde{\gamma_{1}}^{2}}+\frac{1}{a_{2}-\tilde{\gamma_{2}}^{2}})^{3/2}}$, $o_{22}=\frac{4\gamma_{2}^{2}}{(a_{2}-\tilde{\gamma_{2}}^{2})^{3}\sqrt{\frac{1}{a_{1}-\tilde{\gamma_{1}}^{2}}+\frac{1}{a_{2}-\tilde{\gamma_{2}}^{2}}}}-\frac{\gamma_{2}^{2}}{(a_{2}-\tilde{\gamma_{2}}^{2})^{4}(\frac{1}{a_{1}-\tilde{\gamma_{1}}^{2}}+\frac{1}{a_{2}-\tilde{\gamma_{2}}^{2}})^{3/2}}+\frac{1}{(a_{2}-\tilde{\gamma_{2}}^{2})^{2}\sqrt{\frac{1}{a_{1}-\tilde{\gamma_{1}}^{2}}+\frac{1}{a_{2}-\tilde{\gamma_{2}}^{2}}}}$, and is non-negative. The higher dimensions can be generalized with inductions and definitions. Therefore, the overall objective function is convex. Besides, all constraints are linear relations, and thus are convex. Thus, Problem $\mathscr{P}_{2}$ is convex.

### v2

@@ II-B LoRA Dropout
The original dropout technique was proposed in [26] to avoid overfitting during training. In standard dropout, each neuron in the network is dropped from the network with a certain probability. Given that dropout techniques have proven effective in controlling overfitting, we introduce a LoRA dropout framework to enhance generalization when adapting to downstream tasks. The loss function can be re-written as $\mathcal{L}=\frac{1}{|\mathcal{D}|}\sum_{i\in\mathcal{D}}\ell\left(\Delta\bm{\theta}(\bm{m}_{t});\bm{\theta}_{0},{x}_{i}\right),$ where $\Delta{\bm{\theta}}(\bm{m}_{t})$ is the LoRA parameters after the dropout, $\bm{m}_{t}$ is the concatenation of masks of all LoRA modules, which varies across rounds, and $\bm{\theta}_{0}$ is the original parameters of the pre-trained model.
@@ II-B1 Forward Pass
After dropout, the output of the $u$-th layer is given by $\hat{\bm{h}}_{u,i,t}=(\bm{W}_{u,0}+\hat{\bm{B}}_{u,t-1}\hat{\bm{A}}_{u,t-1})\bm{f}_{u-1,i,t}.$
@@ II-C1 Procedure and Algorithm
In this case, each device $k$ holds a local dataset, denoted by $\mathcal{D}_{k}=\{\bm{x}_{i}|i=1,2,\ldots,|\mathcal{D}_{k}|\}$, where $\bm{x}_{i}$ is the $i$-th data sample and the size of $\mathcal{D}_{k}$ is $|\mathcal{D}_{k}|$. Denote $\Delta{\bm{\theta}}(\bm{m}_{k,t})$ as the LoRA parameters after the dropout for the $k$-th device, $\bm{m}_{k,t}$ as the concatenation of all dropout masks of LoRA modules, which varies across different training rounds and different devices, and $\bm{\theta}_{0}$ as the original parameters of the pre-trained model, which has been stored in each device in advance. The objective is to minimize the global loss function, given by $\mathcal{L}=\sum_{k=1}^{K}\tfrac{\left|\mathcal{D}_{k}\right|}{|\mathcal{D}|}\ell_{k}\left(\Delta{\bm{\theta}}(\bm{m}_{k,t});\bm{\theta}_{0},\mathcal{D}_{k}\right),$ where $\mathcal{D}=\{\mathcal{D}_{k}\}$ is the global dataset, the local loss function of the $k$-th device can be written as $\ell_{k}\left(\Delta{\bm{\theta}}(\bm{m}_{k,t});\bm{\theta}_{0},\mathcal{D}_{k}\right)=\tfrac{1}{|\mathcal{D}_{k}|}\sum_{{x}_{i}\in\mathcal{D}_{k}}\ell_{k,i}\left(\Delta{\bm{\theta}}(\bm{m}_{k,t});\bm{\theta}_{0},\bm{x}_{i}\right).$ Similar to (11), FedLoDrop can be written as
@@
  • LoRA Sub-adapter Generation: The server adopts the LoRA dropout technique introduced in Section II-B to generate a sub-adapter for each device. Crucially, the dropout masks are statistically independent across rounds and generated prior to any client’s local computation, thus preventing privacy leaks.
@@ III-B Generalization Error Analysis
In (17), $\ell\left(\bm{\theta},\bm{x}_{j}\right)$ represents the loss for sample $\bm{x}_{j}$ given model parameters $\bm{\theta}$. According to [16], if the following requirements are met: $\ell$ is $\eta$-Lipschitz, $\bm{\theta}_{\ell_{k,\lambda}}(\mathcal{D}_{k})$ is close to $\bm{\theta}_{\ell_{k,\lambda_{t}}}(\mathcal{D}_{k}^{j})$, Hessian matrix $\nabla^{2}{\ell_{k,\lambda_{t}}}(\bm{\theta}_{\ell_{k,\lambda_{t}}}(\mathcal{D}_{k}))$ at $\bm{\theta}_{\ell_{k,\lambda_{t}}}(\mathcal{D}_{k})$ is positive-semidefinite with a singular value decomposition $U_{k,t}\text{diag}(\Lambda_{k,t})U_{k,t}^{-1}$, and $\Lambda_{k,t,\text{min}}=\text{min}\{\Lambda_{k,t,1},\cdots,\Lambda_{k,t,m}\}$, then the LoRA dropout algorithm optimizing $\ell_{k,\lambda_{t}}$ on $|\mathcal{D}_{k}|$ has an upper bound of PHS of  [^2] We only require the regularization coefficient to be sufficiently large (i.e., $\lambda_{t}\geq-\frac{1}{2}\Lambda_{k,t,\min}$), without invoking stronger assumptions about the objective function. Since uniform stability is stricter and argument stability focuses on parameters, we adopt PHS. Despite the non-convexity of LLM training, which may introduce theoretical gaps, the resulting bound provides meaningful qualitative insights into the client numbers and dropout rates. Tighter bounds under weaker assumptions remain an exciting direction for future work.
@@
Theorem 3 elucidates the connection between generalization error and dropout rate. It is shown that increasing the dropout rate decreases the gap between empirical and generalization errors but concurrently increases empirical error. Therefore, an appropriate dropout rate would help balance a trade-off between overfitting and underfitting. By enforcing sparsity more stringently, a larger $\lambda_{t}$ leads to more compact parameter updates and a lower theoretical generalization error, thereby reducing overfitting risk and improving expected out-of-sample performance. While the theorem does not explicitly characterize the relationship between generalization error and LoRA rank, it suggests that decreasing the rank-enhancing parameter sparsity-typically correlates with reduced generalization error. This implies that a model with a lower effective rank tends to have better generalization performance [9].
@@ III-C Upper Bounding Loss Descent
$\quad-\alpha_{B}\Delta\hat{\bm{B}}_{u,t-1}\bm{A}_{u,t-1}$
@@ VI-A Effects of Dropout Rate
FedIT with Gaussian Dropout (0.2) 62.08 77.98 90.20 96.22 94.25 91.72 85.41
FedIT with Gaussian Dropout (0.3) 61.82 77.98 89.22 95.41 93.85 91.47 84.96
@@ VI-D Effects of Parameters and Models
Fig. 6 presents the effects of parameters on the performance. From Fig. 6(a) and Fig. 6(b), one can see that in both models, a larger number of devices generally results in a lower testing accuracy due to the difficulty in adapting to one global model. We further evaluate the sampled setting with 50 devices on LLaMA, where the accuracy decreases modestly to 26.57. As shown in Fig. 6(c), with $\gamma=0.3$, performance first improves with rank but slightly declines when the rank becomes large due to overfitting. For RTE, we further include results under dropout rates of 0.2 and 0.4. It can be observed that insufficient dropout (0.2) leads to overfitting at high ranks, while strong dropout (0.4) can initially cause underfitting but enables higher ranks to be effectively utilized.
@@ A-B PHS upper bound of FedLoDrop
Integrating the result above to Lemma 5 and substituting the regularization coefficient finalizes the proof.
@@ Appendix D Proof Of Lemma 4
where, $o_{11}=\frac{4\gamma_{1}^{2}}{(a_{1}-\tilde{\gamma_{1}}^{2})^{3}\sqrt{\frac{1}{a_{1}-\tilde{\gamma_{1}}^{2}}+\frac{1}{a_{2}-\tilde{\gamma_{2}}^{2}}}}-\frac{\gamma_{1}^{2}}{(a_{1}-\tilde{\gamma_{1}}^{2})^{4}(\frac{1}{a_{1}-\tilde{\gamma_{1}}^{2}}+\frac{1}{a_{2}-\tilde{\gamma_{2}}^{2}})^{3/2}}+\frac{1}{(a_{1}-\tilde{\gamma_{1}}^{2})^{2}\sqrt{\frac{1}{a_{1}-\tilde{\gamma_{1}}^{2}}+\frac{1}{a_{2}-\tilde{\gamma_{2}}^{2}}}}$, $o_{12}=o_{21}=-\frac{\gamma_{1}\gamma_{2}}{(a_{1}-\tilde{\gamma_{1}}^{2})^{2}(a_{2}-\tilde{\gamma_{2}}^{2})^{2}(\frac{1}{a_{1}-\tilde{\gamma_{1}}^{2}}+\frac{1}{a_{2}-\tilde{\gamma_{2}}^{2}})^{3/2}}$, $o_{22}=\frac{4\gamma_{2}^{2}}{(a_{2}-\tilde{\gamma_{2}}^{2})^{3}\sqrt{\frac{1}{a_{1}-\tilde{\gamma_{1}}^{2}}+\frac{1}{a_{2}-\tilde{\gamma_{2}}^{2}}}}-\frac{\gamma_{2}^{2}}{(a_{2}-\tilde{\gamma_{2}}^{2})^{4}(\frac{1}{a_{1}-\tilde{\gamma_{1}}^{2}}+\frac{1}{a_{2}-\tilde{\gamma_{2}}^{2}})^{3/2}}+\frac{1}{(a_{2}-\tilde{\gamma_{2}}^{2})^{2}\sqrt{\frac{1}{a_{1}-\tilde{\gamma_{1}}^{2}}+\frac{1}{a_{2}-\tilde{\gamma_{2}}^{2}}}}$, and is non-negative. The higher dimensions can be generalized with inductions and definitions. Therefore, the overall objective function is convex. Besides, all constraints are linear relations, and thus are convex.